OpenStudy (anonymous):

Simplify the product. (x-4) (x+3)

OpenStudy (lgbasallote):

use FOIL multiply first terms multiply outside terms multiply inside terms multiply last terms

OpenStudy (anonymous):

\[\huge (x + 3)(x - 4)\] \[\huge = x*x + 3*x - x*4 - 3*4\] \[\huge = x^2 + 3x - 4x - 12 \] \[\huge = x^2 - x - 12 \rightarrow answer\]

OpenStudy (anonymous):

thank you @kmstevens10

OpenStudy (anonymous):

\[\Huge THANKYOU \quad A \quad LOT.........\]

OpenStudy (anonymous):

No, thank you :)